Status Report

ISS On-Orbit Status 21 Dec 2002

By SpaceRef Editor
December 21, 2002
Filed under , ,
ISS On-Orbit Status 21 Dec 2002

All ISS systems continue to function nominally, except as noted previously or below.

Around-the-clock sunshine for the station, while winter begins on our Hemisphere. Another typical Saturday off-duty day for the crew today, featuring only a few scheduled activities.

At wake-up at the regular time (1:00 am EST), followed by the usual post-sleep activities of morning inspection, post-sleep hygiene and breakfast, the crew was congratulated on yesterday’s great work on the IATCS (internal active thermal control system) fill, spare Gas Trap fill, and HCOR (high-rate communications outage recorder) installation work.

Then it was time again for the onboard weekly 3-hr. house keeping chores.  [The house cleaning focuses on removal of food waste products, cleaning of compartments with vacuum cleaner, wet cleaning of surfaces with disinfectants and cleaning of fan screens to avoid temperature rises.  Today’s “wet” portion of the “uborka”/clean-up included treatment of the SM dining table surface and the Node/PMA-1 hatch with “Fungistat” disinfectant.  The crew was also requested to use Fungistat for wet cleaning in the FGB: on the hull surface behind panel 205 and on the onboard cable network trunk (BKS) behind panel 108.]

CDR Kenneth Bowersox again transferred excess condensate water from the Lab CCAA air conditioner’s condensate collection tank to a CWC (contingency water container).

Mandatory daily routine servicing tasks were performed by FE-1 Nikolai Budarin (SM SOSh environment control and life support system maintenance) and Bowersox (PCG-STES and ZCG payload status checkup in the Lab).

Budarin and FE-2/SO Don Pettit were scheduled for their weekly PFCs (private family conferences) via S-band.

All crewmembers performed their daily 2.5-hrs. of physical exercise on TVIS, CEVIS, RED and VELO with load trainer.

MCC-H has uplinked the weekly virus definition file update. The NAV (Norton anti-virus) files install automatically when crewmembers log in with their personal user name.

After yesterday’s HCOR installation in the Lab’s AV3 (avionics #3) rack, the recorder was successful activated and checked out by the ground.

The crew’s “job jar” task list today contains CEO ops, PCS (portable computer system) reboot, EXPRESS rack (ER) data cable inspection, LiOH (lithium hydroxide) canister location search, Lab1O4 and Lab1P4 RFCA (rack flow control assembly) reconfiguration,  SSC (station support computer) client spare HDD (hard disk drive) reload, SSC client file server spare HDD reload, and SSC multicast prep work.

Science Update (Expedition Six — 3rd):

Extra-Vehicular Activity Radiation Monitors (EVARM): Continuing.

GASMAP/Pulmonary Function in Flight (PuFF): Continuing.  Next session in January.  

Renal (Kidney) Stone Experiment:  Continuing (crew taking pills).

Human Research Facility/Workstation (HRF WS):  Continuing.

Foot/Ground Reaction Forces During Space Flight (FOOT): Experimenters were delighted to see CDR Bowersox on FOOT doing so well.  All exercises were successful, boding well for the first data take on Christmas Eve.

Pore Formation and Mobility Investigation (PFMI):  Complete for Inc. 5.  On hold pending further MSG problem resolution. Will remain in MSG until further notice.

Space Acceleration Measurement System (SAMS):  SAMS continues to collect acceleration data for vibratory characterization of the microgravity environment.

Microgravity Acceleration Measurement System (MAMS):  MAMS continues measurement of microgravity environment in the quasi -steady regime for general characterization.  

Protein Crystal Growth-Single Locker Thermal Enclosure System (PCG-STES):   PCG STES temperatures are nominal.

Materials ISS Experiment (MISSE):  In progress. Deployed outside. Nominal and collecting data.

Zeolite Crystal Growth (ZCG):  Experimenters appreciate all the extra effort and great work that the crew put into getting ZCG activated.

Crew Earth Observation (CEO):  Excellent images were received of the entire southern Andean ice field.  These  are ideal for comparison with images taken in December in the last two years. Analysis awaits, but experts expect to see change in glacier tongues.  The context/regional images the crew has provided (for areas surrounding detailed images) are important,- facilitating immediate appreciation of the detailed views, and providing independent data on location.
Today’s optional targets for the CEO (crew earth observations program) were Southern Venezuela (mapping swath left of track. Remotest western Amazonia relatively cloudfree.  Crew was to shoot river patterns and any forest clearing. [ISS permits taking advantage of these occasional cloudfree patches to obtain thorough coverage for the first time]), Costa Rica (winter cloud patterns change over central America, when easterly winds speeds increase dramatically [to the point of damaging structures in the populated highlands of Costa Rica]), Mexico City, Mexico (nadir pass; ESC [electronic still camera]), Atlanta, Georgia (nadir pass; ESC), Gulf of St Lawrence (Bay of Fundy to Newfoundland. Cloudfree region between winter storms:  crew was to document the transition season, especially ice on lakes, and tracks of snowfall in the wake of major storms and downwind of water bodies.  On second pass, cloudfree region between winter storms.  Of interest: documenting ice on lakes and snowfall tracks in the wake of storms and downwind of water bodies), Albuquerque, New Mexico (nadir pass; ESC), Jarvis Island (nadir and a touch right of track.  Coral reef mapping project needs detailed views), High Plains, USA (obliques left for change of season views:  ice/no-ice status of lakes, and snowfall tracks of recent storms), Saskatchewan R Swamplands (mapping swath of one of only three North American inland deltas requested.  Looking left for a long swampland [100 km, axis parallel with track] with confused drainage.  Global study of inland deltas is in progress. On second pass, mapping swath of one of only three North American inland deltas requested. Again looking left for a long swampland [100 km, axis parallel with track] with confused drainage]), and High Plains, Canada (oblique views left for change of season scenes: ice/no-ice status of lakes, and snowfall tracks of recent storms).
CEO images can be viewed at the website

U.S. and Russian Segment Status (as of yesterday, 2:37 pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On (32 Amp mode). Vozdukh CO2 scrubber is On.  U.S. CDRA CO2 scrubber is Off. TCCS is operational.  BMP Harmful Impurities unit: Absorbent bed #1 in Purify mode, bed #2 in Purify mode.
  • SM Working Compartment:  Pressure (mmHg) — 752; temperature (deg C) — 25.4; ppO2 (mmHg) — 160.6; ppCO2 (mmHg) — 2.5.
  • SM Transfer Compartment:  Pressure (mmHg) — 761; temperature (deg C) — 19.1.
  • FGB Cabin:  Pressure (mmHg) — 756; temperature (deg C) — 21.0.
  • Node:  Pressure (mmHg) — 749.61; temperature (deg C) — 24.0 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 751.38; temperature (deg C) — 22.0; ppO2 (mmHg) — 157.9; ppCO2 (mmHg) — 5.5;
  • Joint Airlock (Equip. Lock):  Pressure (mmHg) — 751.48, temperature (deg C) — 27.0; shell heater temp (deg C) — 27.1, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1:  Shell heater temp (deg C) — 24.1
  • PMA-2:  Shell heater temp (deg C) — 10.6

(n/a = data not available)

Propulsion System (PS):  

  • Total propellant load available (SM + FGB + Progress) — 3781 kg (8335 lb) [as of 12/19/02].

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  Beta Gimbal Assembly (BGA) 2B & BGA 4B both in Autotrack (sun-tracking) mode.
  • SM batteries:  Batteries #5 and 7 disconnected (suspected failed)  All other batteries (6) are in “Partial Charge”: mode.  
  • FGB batteries:  Batteries #4 was replaced yesterday and is in “Cycle” mode; all other batteries (5) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 is in Standby; PCU-2 is in Standby.

Thermal Control Systems:

  • Air conditioner SKV-1 is On, SKV-2 is Off.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup (new patches loaded on both).
  • EXT-2 is On (primary), EXT-1 is off.
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is On (primary); PL-2 MDM is off (diagnostic
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational. Lane 1 is down (as of 11/14).
  • SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.

Attitude Source:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector — U.S. SIGI-1
  • Attitude — U.S. SIGI-2
  • Angular rates — U.S. RGA-1

Flight Attitude:

  • XPOP 3-CMG Momentum Management (yaw/pitch/roll = 0.52/-8.81/-0.04 deg)

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operating.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally.
  • Ku-band is operating nominally (except for occasionally parking of antenna due to XPOP increasing thermal temperatures).
  • Audio subsystem operating nominally.
  • Video subsystem operating nominally.
  • MCOR (medium-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at Lab PDGF with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings.  MT: latched and mated at WS7, with KA power.  POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is Off; Cupola RWS is Off.

SpaceRef staff editor.